# Bayesian Modular Inference for Copula Models with Potentially Misspecified Marginals

### 12 Mar 2026 v1

@@ I.2 Primitive conditions for Assumption 3
$\Delta^{\gamma}_{n}(\eta,\psi):=\prod_{i=1}^{n}\Delta_{\tilde{a}_{i1}}^{\tilde{b}_{i1}}\cdots\Delta_{\tilde{a}_{id}}^{\tilde{b}_{id}}C(u_{i};\psi),$
@@
Now, the $i$-th term in $\Delta^{\gamma}_{n}(\eta,\psi)$ can be represented as
@@
$\Delta^{\gamma}_{n}(\eta,\psi)=\left\{\prod_{i=1}^{n}\prod_{j=1}^{d}\frac{1-\gamma_{j}}{n+1}\right\}\prod_{i=1}^{n}c\left(w_{i};\psi\right).$
Since the posterior $p_{\mathrm{SMI},\gamma}(\psi,\eta\mid\mathcal{D})$ is as a ratio of the pseudo-likelihood $\Delta^{\gamma}_{n}(\eta,\psi)$, the terms $\left\{\prod_{i=1}^{n}\prod_{j=1}^{d}\frac{1-\gamma_{j}}{n+1}\right\}$, and since loss functions are arbitrary up to an additive scale - and since the loss is defined with logarithms - we abuse notation in what follows and set
$\Delta^{\gamma}_{n}(\eta,\psi):=\prod_{i=1}^{n}c\left(w_{i};\psi\right).$
@@
Appendix II Additional results on the application
@@
Table II.1 reports $\Delta_{\text{Major}}(0.1)$ and $\Delta_{\text{Minor}}(0.1)$, as well as Kendall’s $\tau$ and the parameters of the marginal skew-Gaussian copula for all pairs evaluated at the posterior means, $\widehat{\psi}$, for the conventional posterior, the fully cut posterior, and the optimal SMI posterior. $\Delta_{\text{Major}}(0.1)$, $\Delta_{\text{Minor}}(0.1)$, and Kendall’s $\tau$ are estimated using a large sample from the copula model.
Figures II.1, II.2, and II.3 give quantile dependence plots for all pairs for the SMI, the conventional, and the fully cut posterior respectively.
@@
Table II.1: Bond yields. Kendall’s $\tau$ correlation coefficient, asymmetry in the major and minor diagonals at level $\zeta=0.1$, as well as the parameters of the marginal skew-Gaussian copula for all pairs (rounded to 4 digits) evaluated at the posterior means, $\widehat{\psi}$, for the conventional posterior, the fully cut posterior, $\gamma=(0,0,0)^{\top}$, and the optimal SMI posterior.
Figure II.1: Bond yields. $\rho_{LL}(\zeta)$ (blue), $\rho_{UR}(1-\zeta)$ (orange), $\rho_{LR}(\zeta)$ (green), and $\rho_{UL}(1-\zeta)$ (red) versus $\zeta$ under the SMI posterior mean $\widehat{\psi}$ as well as empirical quantile dependencies (gray) for all pairs.
Figure II.2: Bond yields. $\rho_{LL}(\zeta)$ (blue), $\rho_{UR}(1-\zeta)$ (orange), $\rho_{LR}(\zeta)$ (green), and $\rho_{UL}(1-\zeta)$ (red) versus $\zeta$ under the conventional posterior mean $\widehat{\psi}$ as well as empirical quantile dependencies (gray) for all pairs.
Figure II.3: Bond yields. $\rho_{LL}(\zeta)$ (blue), $\rho_{UR}(1-\zeta)$ (orange), $\rho_{LR}(\zeta)$ (green), and $\rho_{UL}(1-\zeta)$ (red) versus $\zeta$ under the fully cut posterior mean $\widehat{\psi}$ as well as empirical quantile dependencies (gray) for all pairs.

### 02 Apr 2026 v2

@@ B.2 Primitive conditions for Assumption 3
$\Delta_{n}(\eta,\psi;\gamma):=\prod_{i=1}^{n}\Delta_{\tilde{a}_{i1}}^{\tilde{b}_{i1}}\cdots\Delta_{\tilde{a}_{id}}^{\tilde{b}_{id}}C(u_{i};\psi),$
@@
Now, the $i$-th term in $\Delta_{n}(\eta,\psi;\gamma)$ can be represented as
@@
$\Delta_{n}(\eta,\psi;\gamma)=\left\{\prod_{i=1}^{n}\prod_{j=1}^{d}\frac{1-\gamma_{j}}{n+1}\right\}\prod_{i=1}^{n}c\left(w_{i};\psi\right).$
Since the posterior $p_{\mathrm{SMI},\gamma}(\psi,\eta\mid\mathcal{D})$ is as a ratio of the pseudo-likelihood $\Delta_{n}(\eta,\psi;\gamma)$, the terms $\left\{\prod_{i=1}^{n}\prod_{j=1}^{d}\frac{1-\gamma_{j}}{n+1}\right\}$, and since loss functions are arbitrary up to an additive scale - and since the loss is defined with logarithms - we abuse notation in what follows and set
$\Delta_{n}(\eta,\psi;\gamma):=\prod_{i=1}^{n}c\left(w_{i};\psi\right).$
@@
Appendix C Additional information for the simulation study
We consider the three evaluation metrics
Appendix D Additional information on the bond yields application
D.1 Parameterization of the skew-normal copula
Assume $\widetilde{Y}_{t}=(Y_{\text{VIX},t},Y_{\text{AAA},t},Y_{\text{BBB},t})^{\top}$ follows a stationary stochastic process. Then from Theorem 1 in Smith (2015), the three marginals of $\widetilde{Y}_{t}$ are time invariant, and the copula of $Y_{t}=(\widetilde{Y}_{t}^{\top},\widetilde{Y}_{t-1}^{\top})^{\top}$ is also stationary, in that $C(u_{t},u_{t-1})$ is invariant to $t$. Invariance in the implicit SN copula $C$ follows from invariance in the underlying SN distribution, which can be enforced by the two parameter constraints described below.
$\Omega=\begin{pmatrix}\Omega_{0}&\Omega_{1}\\ \Omega_{1}^{\top}&\Omega_{0}\end{pmatrix}\,$. (D.12)
The second constraint follows from writing $\alpha=\left(1-\delta^{\top}\Omega^{-1}\delta\right)^{-\frac{1}{2}}\Omega^{-1}\delta$ in terms of the slant parameters $\delta$ in Azzalini and Valle (1996). If we partition $\delta=(\delta_{2}^{\top},\delta_{1}^{\top})^{\top}$ and $\alpha=(\alpha_{2}^{\top},\alpha_{1}^{\top})^{\top}$ consistent with (D.12), then invariance follows when $\delta_{2}=\delta_{1}$. This implies that $\alpha_{1}=(\Omega_{0}-\Omega_{1})^{-1}(\Omega_{0}-\Omega_{1}^{\top})\alpha_{2}$. All together, $\Omega_{0}$ is parameterized through 3 unique transformed partials, $\Omega_{1}$ is parameterized through 9 unique transformed partials, and $\alpha$ is parameterized by 3 unique values in $\alpha_{2}$, giving an unrestricted copula parameter vector $\psi$ of dimension 15. Because there are only three unique marginals, $\gamma$ is of dimension three.
D.2 Additional results
@@
Figure D.1 shows how the utility improves with the number of SMI posteriors checked. In this example, BO converges fast after a brief exploration of the $\gamma$ space.
Table D.1 reports $\Delta_{\text{Major}}(0.1)$ and $\Delta_{\text{Minor}}(0.1)$, as well as Kendall’s $\tau$ and the parameters of the marginal skew-Gaussian copula for all pairs evaluated at the posterior means, $\widehat{\psi}$, for the conventional posterior, the fully cut posterior, and the optimal SMI posterior. $\Delta_{\text{Major}}(0.1)$, $\Delta_{\text{Minor}}(0.1)$, and Kendall’s $\tau$ are estimated using a large sample from the copula model.
Figures D.2, D.3, and D.4 give quantile dependence plots for all pairs for the SMI, the conventional, and the fully cut posterior respectively.
Figure D.1: Convergence of BO over $100$ iterations for the bond yields example. The vertical axis is the utility $u(\gamma)$, and the black line indicates the optimal value $u(\gamma^{\ast})$ against number of $\gamma$ values explored. The values $u((1,1,1)^{\top})$, $u((0,0,0)^{\top})$ are plotted as gray horizontal lines.
@@
Table D.1: Bond yields. Kendall’s $\tau$ correlation coefficient, asymmetry in the major and minor diagonals at level $\zeta=0.1$, as well as the parameters of the marginal skew-Gaussian copula for all pairs (rounded to 4 digits) evaluated at the posterior means, $\widehat{\psi}$, for the conventional posterior, the fully cut posterior, $\gamma=(0,0,0)^{\top}$, and the optimal SMI posterior.
Figure D.2: Bond yields. $\rho_{LL}(\zeta)$ (blue), $\rho_{UR}(1-\zeta)$ (orange), $\rho_{LR}(\zeta)$ (green), and $\rho_{UL}(1-\zeta)$ (red) versus $\zeta$ under the SMI posterior mean $\widehat{\psi}$ as well as empirical quantile dependencies (gray) for all pairs.
Figure D.3: Bond yields. $\rho_{LL}(\zeta)$ (blue), $\rho_{UR}(1-\zeta)$ (orange), $\rho_{LR}(\zeta)$ (green), and $\rho_{UL}(1-\zeta)$ (red) versus $\zeta$ under the conventional posterior mean $\widehat{\psi}$ as well as empirical quantile dependencies (gray) for all pairs.
Figure D.4: Bond yields. $\rho_{LL}(\zeta)$ (blue), $\rho_{UR}(1-\zeta)$ (orange), $\rho_{LR}(\zeta)$ (green), and $\rho_{UL}(1-\zeta)$ (red) versus $\zeta$ under the fully cut posterior mean $\widehat{\psi}$ as well as empirical quantile dependencies (gray) for all pairs.